Write a balanced reaction for which the following rate relationships are true:
Rate = -1/2 delta[N2O5] over delta t = 1/4 delta[NO2] over delta t = delta]O2] over delta t

a. 2N2O5 yields 4NO2 + O2
b. 4NO2 + O2 yields 2N2O5
c. 2N2O5 yields NO2 + 4O2
d. 1/4 NO2 + O2 yields 1/2 N2)5
e. 1/2 N2O5 yields 1/4 NO2 + O2

Can someone please describe WHY the answer is A? Please show work.
Answer #1

a. 2N2O5 yields 4NO2 + O2  

2N2O5 -------> 4NO2 + O2

-1/2\Delta[N2O5]/\Deltat   = +1/4\Delta[NO2]/\Deltat   = +\Delta[O2]/\Deltat

Ex xA + yB ------> zC

-1/x\Delta[A]/\Deltat = -1/y \Delta[B]/\Deltat = +1/z\Delta[C]/\Deltat